# Conformal Prediction with Upper and Lower Bound Models

### v1

@@ 5 Experiments
Furthermore, CPUL and CPUL-OMLT consistently achieve state-of-the-art performance, across datasets and target coverage. On the other hand, CQR and CQR-r exhibit the worst performance overall, likely due to their inability to properly address heteroskedasticity. This is most evident in Figure 3. On the 1354_pegase dataset (which corresponds to a larger power grid) CPUL-OMLT achieves the smallest interval size for both $\alpha=5\%$ and $\alpha=10\%$. In particular, CPUL-OMLT produces prediction intervals whose witdth is about 10% smaller than the second-best method, CPUL. These results highlight the important of the model selection procedure in CPUL.
@@
Figure 3 demonstrates that, while methods like Split CP and SFD-CP deliver good results occasionally, their performance is highly variable across different datasets and confidence levels. For instance, although SFD-CP is among the best-performing methods on 89_pegase, its performance on 1354_pegase is significantly worse than CPUL when coverage is close to 100% (corresponding to small values of $\alpha$). Similarly, Split CP using $\hat{B}^{u}$ is among the worst performers on the 118_IEEE dataset, but among the top performers on the 1354_pegase dataset. Such variability in performance further reinforce the value of model selection, which is most evident on the 11354_pegase dataset, Thereby, methods such as SCP with $\hat{B}^{l}$ and SFP CP perform well when the coverage is between 85% and 90%, but are out-performed by SCP with $\hat{B}^{u}$ when coverage is above 90%. In contrast, CPUL and CPUL-OMLT perform well across the entire coverage range. This is explained by the fact that CPUL and CPUL-OMLT select the best construction for each target coverage $\alpha$.
@@
Finally, while CPUL and CPUL-OMLT perform similarly on the 89_pegase and 118_IEEE datasets, CPUL-OMLT offers a clear improvement on the 1354_pegase dataset, especially for coverage levels above 90% (see Figure 3). A more granular analysis of each model’s behavior reveals that CPUL’s performace is worse on samples where $[\hat{B}^{l},\hat{B}^{u}]$ is small. Namely, CPUL achieves a coverage of only about 50% across the 5% of test samples with smallest initial interval $[\hat{B}^{l},\hat{B}^{u}]$. Recall that these samples correspond to regions where the valid lower and upper bounds $\hat{B}^{l},\hat{B}^{u}$ are the most accurate, which demonstrates the issue of paradoxical miscoverage (see Section 4 and Figure 2). CPUL-OMLT effectively mitigates this issue, which results in better coverage and smaller interval length overall.
@@ 6 Conclusion
This paper introduced CPUL, a novel CP mechanism designed for the settings that determinists lower and upper bounds on the target variable are available. By integrating multiple interval constructions strategies within nested CP framework, CPUL effectively leverages the structure of these bounds to improve efficiency. The paper also introduces the OMLT mechanism, which leverages the strong condition that tight initial bounds already provide highly precise confidence intervals. Failure to account for this often leads to undercoverage in such regions, a challenge that existing methods struggle to address. Experimental validations were conducted on optimization problems across three power systems datasets, demonstrating that the proposed approach consistently outperforms the traditional CP methods by providing better efficiency. The proposed CPUL-OMLT method provides state-of-the-art efficiency (i.e., low average interval length) and remains consistent across different datasets, highlighting its robustness and practical relevance.
While this paper prioritizes achieving simpler marginal coverage with minimal computational overhead, future research could extend the framework to enforce stronger conditional coverage guarantees, leading to more adaptive and precise intervals. Exploring extensions to high-dimensional and structured data could further broaden the applicability of CPUL-OMLT in complex real-world settings, such as energy systems and beyond.
ACKNOWLEDGMENT
This research was partly funded by NSF award 2112533, ARPA-E PERFORM award AR0001136, and NSF award DGE-2039655. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of NSF or ARPA-E.
@@ A.1 Problem formulations
Consider a parametric optimization problem with parameter $x\in\mathcal{X}$, of the form
$\Phi(x)=\min_{s}\quad$ $f_{x}(s)$ (30a)
s.t. $g_{x}(s)\leq 0$ (30b)
$h_{x}(s)=0$ (30c)
where $s\in\mathbb{R}^{n}$ denotes decision variables, $f_{x}:\mathbb{R}^{n}\mapsto\mathbb{R}$ is the objective function, and $g_{x}:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ and $h_{x}:\mathbb{R}^{n}\mapsto\mathbb{R}^{p}$ are the constraints. The optimal value of (30) is $\Phi(x)$, and $\Phi$ is called the value function. Denote by $\mathbf{X}(x)$ the feasible set of (30), i.e,
$\mathbf{X}(x)=\{s\in\mathbb{R}^{n}|g_{x}(s)\leq 0,h_{x}(s)=0\}$
A candidate solution $s\in\mathbb{R}^{n}$ is primal feasible if $s\in\mathbf{X}(x)$, i.e., if $g_{x}(s)\leq 0$ and $h_{x}(s)=0$, otherwise it is infeasible. Suppose that there exists a unique optimal solution $x*$. Then, $f_{x}(x*)\leq f_{x}(s)$ for all $s\in\mathbf{X}(x)$
@@ A.2 Data generation
Note that all CP methods require true labels to compute exact residuals during the calibration step (12), and the test set must be labeled as well for performance evaluation. These true labels are computed using the LP solver HiGHS highs, via OPFGenerator opf_generator. For the significantly larger training set $\mathcal{D}_{\mathrm{train}}$, true labels are not required, as the training process is self-supervised.
The three datasets used in this study are based on selected snapshots from the Power Grid Lib - Optimal Power Flow collection pglib.The sampling distribution is consistent across all cases. For each sample $i$, each load $d_{l}$ is generated as
@@ A.3.1 Primal feasible proxies
A primal-feasible solution to 33 can be obtained by following the procedure below, similar to chen2023end:
@@ A.3.2 Dual feasible proxies
The Dual Lagrangian Learning framework tanneau2024dual is applied;
the specific dual recovery procedure reads as follows:
@@ A.4.1 Details of comparison methods
CPUL and its optimized version, CPUL-OMLT, are compared against three other CP methods. For clarity, the construction of all methods used in the experiments is summarized in Table 1.

### v2

@@ 5 Experiments
Furthermore, CPUL and CPUL-OMLT consistently achieve state-of-the-art performance, across datasets and target coverage. On the other hand, CQR and CQR-r exhibit the worst performance overall, likely due to their inability to properly address heteroskedasticity. This is most evident in Figure 2. On the 1354_pegase dataset (which corresponds to a larger power grid) CPUL-OMLT achieves the smallest interval size. In particular, CPUL-OMLT produces prediction intervals whose witdth is about 10% smaller than the second-best method, CPUL. These results highlight the important of the model selection procedure in CPUL.
@@
Figure 2 demonstrates that, while methods like Split CP and SFD-CP deliver good results occasionally, their performance is highly variable across different datasets and confidence levels. For instance, although SFD-CP is among the best-performing methods on 89_pegase, its performance on 1354_pegase is significantly worse than CPUL when coverage is close to 100% (corresponding to small values of $\alpha$). Similarly, Split CP using $\hat{B}^{u}$ is among the worst performers on the 118_IEEE dataset, but among the top performers on the 1354_pegase dataset. Such variability in performance further reinforce the value of model selection, which is most evident on the 1354_pegase dataset, Thereby, methods such as SCP with $\hat{B}^{l}$ and SFP CP perform well when the coverage is between 85% and 90%, but are out-performed by SCP with $\hat{B}^{u}$ when coverage is above 90%. In contrast, CPUL and CPUL-OMLT perform well across the entire coverage range. This is explained by the fact that CPUL and CPUL-OMLT select the best construction for each target coverage $\alpha$.
@@
Finally, while CPUL and CPUL-OMLT perform similarly on the 89_pegase and 118_IEEE datasets, CPUL-OMLT offers a clear improvement on the 1354_pegase dataset, especially for coverage levels above 90% (see Figure 2). A more granular analysis of each model’s behavior reveals that CPUL’s performace is worse on samples where $[\hat{B}^{l},\hat{B}^{u}]$ is small. Namely, CPUL achieves a coverage of only about 50% across the 5% of test samples with smallest initial interval $[\hat{B}^{l},\hat{B}^{u}]$. Recall that these samples correspond to regions where the given bounds $\hat{B}^{l},\hat{B}^{u}$ are the most accurate, which demonstrates the issue of paradoxical miscoverage (see Section 4 and Figure 1(b)). CPUL-OMLT effectively mitigates this issue, which results in better coverage and smaller interval length overall.
@@ 6 Conclusion
This paper introduced CPUL-OMLT, a novel CP mechanism designed for the settings that deterministic lower and upper bounds on the target variable are available. By integrating multiple interval construction strategies within the NCP framework, CPUL effectively leverages the structure of these bounds to improve efficiency. The OMLT mechanism leverages the strong condition that tight initial bounds already provide highly precise confidence intervals. Failure to account for this often leads to undercoverage in such regions, a challenge that existing methods struggle to address. Experimental validations were conducted on optimization problems across three power systems datasets, demonstrating that the proposed approach consistently outperforms the traditional CP methods by providing better efficiency. The proposed CPUL-OMLT method provides state-of-the-art efficiency and remains consistent across different datasets, highlighting its robustness and practical relevance.
Future work could extend CPUL-OMLT beyond the efficient marginal coverage studied here. Potential directions include enforcing stronger conditional coverage to incorporate the already computed bounds $\hat{B}^{l},\hat{B}^{u}$, integrating with optimization algorithms like branch and bound to directly boost downstream task performance, and adapting the method for high-dimensional/structured data to increase applicability in complex settings like energy systems.
@@ A.1 Problem formulations
With a slight abuse of notation for the readability, here equality constraints are written out explicitly. Let $\mathbf{S}(x)$ denote the feasible set of, i.e,
$\mathbf{S}(x)=\{s\in\mathbb{R}^{n}|g_{x}(s)\leq 0,h_{x}(s)=0\}$
A candidate solution $s\in\mathbb{R}^{n}$ is primal feasible if $s\in\mathbf{S}(x)$, i.e., if $g_{x}(s)\leq 0$ and $h_{x}(s)=0$, otherwise it is infeasible. Suppose that there exists a unique optimal solution $x^{*}$. Then, $f_{x}(x^{*})\leq f_{x}(s)$ for all $s\in\mathbf{S}(x)$.
@@ A.2 Data generation
Note that all CP methods require true labels to compute exact residuals during the calibration step (11), and the test set must be labeled as well for performance evaluation. These true labels are computed using the LP solver HiGHS highs, via PGLearn opf_generator. For the significantly larger training set $\mathcal{D}_{\mathrm{train}}$, true labels are not required, as the training process is self-supervised. The three datasets used in this study are based on selected snapshots from the Power Grid Lib - Optimal Power Flow collection pglib. The sampling distribution is consistent across all cases. For each sample $i$, each load $d_{l}$ is generated as
@@ A.3.1 Primal feasible proxies
A primal-feasible solution to 29 can be obtained by following the procedure below, similar to chen2023end:
@@ A.3.2 Dual feasible proxies
The Dual Lagrangian Learning framework tanneau2024dual is applied; the specific dual recovery procedure reads as follows:
@@ A.4.1 Details of comparison methods
CPUL and its optimized version, CPUL-OMLT, are compared against three other CP methods. For clarity, the construction of all methods used in the experiments is summarized in Table LABEL:table:baseline.